OpenStudy (gabylovesyou):

Find the diameter of a cone that has a volume of 56.52 cubic inches and a height of 6 inches. Use 3.14 for pi. 3 inches 6 inches 9 inches 12 inches

OpenStudy (gabylovesyou):

3 ?

OpenStudy (anonymous):

a. 3inches 56.52/3.14=18 18/6=3 3*(3/1)= 9 9square rooted = 3

OpenStudy (anonymous):

ya

OpenStudy (gabylovesyou):

lol ok ty <3 ;D